OpenStudy (czesc):

INSTANT FAN / MEDAL Find the area of the region that is inside the graph of r = 1 + sin θ and outside of the graph of r = 1.

OpenStudy (czesc):

1.785 5.571 2.785 1.393 ARE possible answers

OpenStudy (anonymous):

First equal two equations to find intersections and in which angles that happens: \(1+\sin{\theta }=1\) \( \theta = 0 \),\( \theta = 2\pi \)

OpenStudy (anonymous):

now, calc the integral: \[ A=\int_{0}^{2\pi} \frac{1}{2} [(1+\sin{\theta })^2-(1)^2]d\theta \]

OpenStudy (czesc):

@M4thM1nd when I plug that into my calculator I get -1.57//-pi/2

OpenStudy (czesc):

Yeah that's getting a negative answer.

OpenStudy (czesc):

what'd you guys get for the area?